OpenStudy (anonymous):

y = -1/3x +3 -x + 3y = 9 Type the first equation from the group you selected and identify the slope and the y-intercept of the equation. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. Type the second equation from the group you selected and change the equation into slope-intercept form. Be sure to show all of your work. Identify the slope and y-intercept of the equation. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. In a paragraph of at least three se

OpenStudy (blurbendy):

slope for #1 = -1/3, y-intercept = 3 slope for #2 = 1/3, y-intercept = 3

OpenStudy (anonymous):

Can you show your work?

OpenStudy (anonymous):

@saifoo.khan or @phi could you help? Pretty please?

OpenStudy (anonymous):

I havent done this in a longg time... but this might help? :) http://answers.yahoo.com/question/index?qid=20110317163112AA02c0x

OpenStudy (anonymous):

@shadowfiend Help please?

OpenStudy (blurbendy):

For the first equation, y = -1/3x + 3, this is in the form y = mx + b. Your slope is always the "m" value, so that is why the slope is -1/3. The y-intercept is the +b. That is y the y-intercept is 3

OpenStudy (blurbendy):

for the second equation, it's not in the form y = mx +b. So to get it in that form, you first add x to both sides to get 3y = x +9. Almost there. Divide both side by 3. y = 1/3x + 9. So, the slope is 1/3, and the y-intercept is 9. Understand?